Leah puts $3,000 into her savings account that earns 1.8% interest compounded quarterly.
a) What will the balance be after 5 years? Round to the nearest cent.
b) How many years will it take her to reach a balance of $4,000? Round to three decimal places.
Can you explain me this second homework question step-by step please? Thank you!

Answer by mananth(16946)


Leah puts $3,000 into her savings account that earns 1.8% interest compounded quarterly.
P=3000
r=1.8% = 0.018
n=4
t=5
A = 3000(1+0.018/4) ^ 4*5
A= 3000*(1+0.0045)^20
the balance be after 5 years $3281.86
b) How many years will it take her to reach a balance of $4,000?
4000 = 3000*(1+0.0045)^4t
4/3 = 1.0045^4t
log(4/3) = 4t*log(1.0045)
(log(4/3))/(4* log (1.0045))=t
t=16.018 years